OpenStudy (anonymous):

Medal A car and a cyclist start at 10 am from the same point, headed in the same direction. The average speed of the car is 5 mph more than three times the average speed of the cyclist. In 1.5 h, the car is 46.5 mi ahead of the cyclist. Find the rate of the cyclist.

OpenStudy (anonymous):

Let x be the average speed of the cyclist. Since the average speed of the car is 5mph more than 3 times the cyclist's, then his speed is 3x + 5. Since they are moving in the same direction, the distance between them will increase by the differences of their speed. For example, if a car is moving 50mph and another card is moving 40mph, then the distance between two cars increase at 50 - 40 = 10mph. Since the distance is 46.5 miles, we get: 46.5 = 1.5(3x + 5) - 1.5x 46.5 = 4.5x + 7.5 - 1.5x 46.5 = 3x + 7.5 39 = 3x x = 13 Since x is the speed of the cyclist, the speed of the cyclist is 13mph.

OpenStudy (anonymous):

ohhh i c thank you soo much :)